OpenStudy (aristasiebrits8):

Hello Everybody!! I really need help on this Algebra question. I have been working on my Math all day, and now I can't think clear anymore. I just need help on this one question, and I will MEDAL and FAN anybody that helps me answer this. A system of linear inequalities is shown below: x + y ≥ 4 y < 2x - 3 Describe the steps to graph the solution set to the system of inequalities. Thank you very much!

OpenStudy (aristasiebrits8):

@phi @pooja195 Please help!

zepdrix (zepdrix):

If you take the first equation, and subtract x from each side, then here is our system,\[\large\rm y\ge-x+4\]\[\large\rm y<2x-3\] Start out this process by graphing these two lines,\[\large\rm y=-x+4\]\[\large\rm y=2x-3\]Do you understand how to do that?

OpenStudy (rayep12):

The first thing I tell my students about these is to first temporarily replace the inequality signs with = signs, simply for the purposes of solving for y and graphing the lines. We will put them back in later. For the first one, we will rename that x + y = 4. Solving this for 4 gives y = 4 - x. Are you familiar with the slope-intercept form of graphing? The y-intercept is 4 and the slope is negative 1.

OpenStudy (aristasiebrits8):

Yes. I will show you in a second.

OpenStudy (aristasiebrits8):

Here it is. Is it correct?

OpenStudy (aristasiebrits8):

@zepdrix are you still there?

zepdrix (zepdrix):

Ummm first of all... you used a calculator... I don't see how that is "your attempt". And besides that.. the numbers don't match... I'm so confused... why do you have .037x when our equation says 2x?? ?? 0_o

OpenStudy (aristasiebrits8):

I didn't use a calculator, I used desmos. And I don't know I'll check again. Sorry

OpenStudy (aristasiebrits8):

Sorry! I had two files under the same name...

zepdrix (zepdrix):

:P

OpenStudy (aristasiebrits8):

Let's try again...

OpenStudy (aristasiebrits8):

@zepdrix is it better?

zepdrix (zepdrix):

Oh, I'm not willing to download a docx >.< no desmos link or picture? :3

OpenStudy (aristasiebrits8):

OKay. Coming

OpenStudy (aristasiebrits8):

Okay! Here it is...

zepdrix (zepdrix):

i wish you weren't using a graphing calculator to do this :P but whatever... first equation has a starting point of 4